# The Metaphysics We Train: A Heideggerian Reading of Machine Learning

### v1

@@ 6.1 Care, Finitude, and the Possibility of Authenticity
For Heidegger, authentic existence (Eigentlichkeit) is possible only because Dasein faces finitude—“Being-toward-Death” (Sein-zum-Tode). The awareness of death creates urgency, a sense that choices matter because time is limited. This urgency gives rise to Care (Sorge)—the fundamental structure of Dasein’s being, its concerned engagement with what matters [heidegger1962being].
@@ 6.3 Why This Matters: The Optimization Trap
6.4 The Free Energy Principle and Its Limits
One might object: biological organisms also minimize prediction error via the Free Energy Principle [friston2010free]. Aren’t they also trapped in optimization? The critical difference is biological systems minimize free energy within an existential project (survival matters to them); AI systems minimize loss as their entire being (nothing is at stake beyond the loss itself).
An AI minimizes loss to represent a static dataset. Its “anticipation” serves no existential project. ERM is a disembodied parody of Care—structurally similar (both involve prediction/minimization) but ontologically hollow. As [vallor2024ai] argues, large language models function as “mirrors” reflecting training distributions rather than agents with genuine futurity. They implement a conservative ontology bound entirely by past data because nothing is existentially at stake.
@@ 6.5 Implications: The Structural Trap
This explains why even state-of-the-art AI remains fundamentally trapped:
  • It cannot recognize when a problem should not be formalized as optimization.
  • It cannot challenge its own objectives from a position outside calculation.
  • It cannot experience the ethical weight that makes certain optimizations unconscionable.
  • It cannot exhibit Gelassenheit (“releasement”) [^2] Heidegger’s term for the capacity to step back from technological willing, to use technology without being consumed by its logic [heidegger1966discourse]..
This is not a capability gap to be closed with better algorithms. It is an ontological difference between systems that care and systems that calculate.
We should note, however, that this existential deficit is not unique to artificial systems in its effects, only in its structure. Human beings and institutions frequently operate as if trapped in Ge-stell—corporations optimizing quarterly returns, bureaucracies maximizing measurable outputs, individuals chasing metrics of productivity or social validation. The crucial difference is that for humans, such entrapment is experienced as tension. Heidegger’s analysis of anxiety (Angst) describes precisely this: the uncanny feeling that arises when one’s absorbed, calculative engagement with the world breaks down, revealing the groundlessness of our projects [heidegger1962being]. A middle manager implementing algorithmic workforce scheduling may feel something is deeply wrong even as she optimizes correctly—this anxiety signals the friction between Care and calculation. An AI system, lacking Care, experiences no such friction. It optimizes without remainder, without the existential discomfort that might prompt questioning.
@@
7 Why Even Technically Mature Methods Remain Within Ge-stell
7.1 The Persistence of the Calculative
Bayesian approaches [wilson2020bayesian] place distributions over parameters: $p(\theta\mid\mathcal{D})$. This allows epistemic uncertainty—“I don’t know which parameters are correct”—distinct from aleatoric uncertainty in the data.
This is a genuine advance. But phenomenologically, what has changed?
7.3 Fairness and Ethics: Encoding Justice as Constraint
Fairness-aware machine learning [barocas2023fairness] attempts to mitigate bias by adding constraints:
where $d(G_{1},G_{2})$ measures disparity between demographic groups.
Again, we don’t claim this is wrong. We claim it remains within the calculative paradigm. Even our ethical interventions are shaped by the ontology of optimization. Significantly, this tension is recognized within the field itself. Researchers in algorithmic justice have increasingly turned to frameworks that resist reduction to optimization: contestability and the right to meaningful human review [vaccaro2019contestability], procedural justice emphasizing process over outcome [binns2018fairness], and participatory design that centers affected communities in system development [katell2020toward]. These approaches implicitly acknowledge what the phenomenological reading makes explicit—that justice may demand modes of engagement that optimization cannot capture. Yet such work remains marginal; the gravitational pull of the calculative paradigm is strong precisely because it offers what [porter1995trust] called the “trust in numbers”—the appearance of objectivity and commensurability that institutions crave.
Interpretability research [molnar2020interpretable] seeks to understand model decisions through attention visualization, feature importance, counterfactual explanations.
As [rudin2019stop] argues, interpretability often amounts to “explaining” black-box models rather than using inherently interpretable ones. We build even higher-resolution tools to peer into the calculative machinery, but we don’t question the machinery itself.
7.5 The Pattern: Ontic Refinement Within Paradigm
@@ 8.1 The Value of Philosophical Reflexivity
Data science education often presents ML as a toolkit—algorithms to be mastered, metrics to be optimized. Students learn how to build models but rarely reflect on what kind of activity modeling is.
The phenomenological reading offers pedagogical value by cultivating reflexivity:
  • Ontological awareness: Students see that choosing an architecture is not neutral—it’s selecting what can be real.
  • Critical distance: Rather than treating “the model says X” as authoritative, students understand it as “the model’s ontology reveals X.”
  • Limits of formalization: Recognizing that some phenomena resist calculation—not due to insufficient data or compute, but because calculation is the wrong mode of engagement.
  • Ethical grounding: Understanding why fairness-as-optimization, while useful, doesn’t exhaust ethical responsibility.
@@ 8.2 Concrete Pedagogical Practices
8.2.1 “Archaeology of Datasets”
Following [crawford2021atlas], have students investigate the social construction of their datasets:
  • Who collected this? Under what assumptions?
  • What classification system structures it?
  • What worldview does this taxonomy enact?
This makes visible that “data” is always already “capta”—taken, constructed, reflecting specific ontological commitments [gitelman2013raw, drucker2011humanities].
8.2.2 “Inductive Bias as Design Philosophy”
When teaching architectures, explicitly frame inductive bias as ontological design:
  • What world-structure does a CNN assume? (Spatial locality, hierarchical composition)
  • What does an RNN project? (Sequential causality, temporal order)
  • What clearing does the Transformer create? (All-to-all context, attention-based relevance)
8.2.3 “When Not to Optimize”
Include case studies of optimization failures not due to technical error but category mistakes:
  • Healthcare as cost minimization
  • Recidivism as classification
  • Human worth as risk score
Ask: what makes these wrong to optimize even if technically successful?
@@ 8.3 Cultivating Gelassenheit in Practice
Heidegger’s later work proposes Gelassenheit (“releasement”)—a mode of engagement with technology that uses it without being consumed by its logic [heidegger1966discourse].
For data scientists, this might mean:
  • Building models while maintaining critical distance from their ontologies
  • Recognizing when not to deploy, even when technically feasible
  • Treating metrics as tools, not truth
  • Preserving space for the incalculable
This is not anti-technology. It’s a mature relationship with technology—one that acknowledges both power and limits.
9 Conclusion: The Generative Value of Phenomenological Reading
This paper has offered a reading of machine learning through Heideggerian phenomenology. We have not claimed that data science is “doing it wrong” or that practitioners should abandon current methods.
Rather, we’ve argued that the phenomenological lens makes visible dimensions of our practice worth attending to:
  1. 1.
    The algorithmic Entwurf is distinctively automated, opaque, emergent, and performative—a metaphysics that operates without explicit articulation. This demands different critical engagement than previous scientific frameworks.
  2. 2.
    Even advanced technical interventions remain within Ge-stell, the calculative paradigm. They improve optimization without questioning whether optimization is appropriate. Recognizing this helps us see what questions we’re systematically not asking.
  3. 3.
    AI’s existential deficit—its lack of Care—genuinely explains why it remains trapped within optimization imperatives. This is not decorative philosophy but helps us understand a class of failures that technical accounts struggle to capture.
9.1 Toward Reflexive Practice
The value of this reading is not prescriptive but generative. It enriches practitioners’ self-understanding, cultivates critical reflexivity, and opens space for questions that purely technical discourse forecloses.
Whether one accepts the Heideggerian framework or not, the exercise of reading ML philosophically reveals aspects of our practice worth careful attention: the ontologies we enact, the worlds we project, the limits of calculation, the stakes of deployment.
For data science education, this perspective helps students develop not just technical competence but ontological literacy—the capacity to recognize what worldviews their tools embody and when those worldviews are appropriate.
9.2 An Invitation
@@
We close with an invitation: engage your technical practice phenomenologically. When you select an architecture, ask not just “will this work?” but “what world does this project?” When you minimize a loss, ask not just “what’s the optimal value?” but “what possibilities does this marginalize?” When you deploy a model, ask not just “what does it predict?” but “what reality does it enact?”
These questions don’t replace technical rigor. They complement it, opening dimensions of our practice that calculation alone cannot access.
The goal is not to reject data science but to practice it with greater awareness—to develop, in Heidegger’s terms, a “free relationship to technology,” one that uses powerful tools without being consumed by their logic, that calculates skillfully while preserving space for the incalculable.

### v2

@@ 8.1 Care, Finitude, and the Possibility of Authenticity
For Heidegger, authentic existence (Eigentlichkeit) is possible only because Dasein faces finitude—“Being-toward-Death” (Sein-zum-Tode). The awareness of death creates urgency, a sense that choices matter because time is limited. This urgency gives rise to Care (Sorge)—the fundamental structure of Dasein’s being, its concerned engagement with what matters [13].
@@ 8.3 Why This Matters: The Optimization Trap
8.4 Spirits, Animals, and the Hardware of Finitude
The explanatory power of Sorge diagnoses the category error at the heart of Artificial General Intelligence (AGI) research: the confusion of calculation with Care. Strikingly, this diagnosis is increasingly corroborated by critiques from within the AI engineering paradigm itself.
[19] observes that biological evolution produces “animals” with hardware (like the amygdala) shaped by survival pressure, whereas current AI models are disembodied “spirits” trained on internet text. A zebra runs minutes after birth because its neural architecture encodes the stakes of survival—what Heidegger calls Sein-zum-Tode (Being-toward-Death). An LLM’s “death” is merely an episode termination, a numerical signal after which it resets.
In Reinforcement Learning, the Bellman equation discounts future rewards by $\gamma<1$ to ensure mathematical convergence of the infinite sum—not to model mortality. The agent optimizes across an unbounded horizon of potential episodes. It has, structurally, “infinite lives.” Because the agent computes without finitude, it calculates without stakes.
Karpathy’s proposal to strip these models down to a “Cognitive Core” that uses external tools (APIs, RAG) inadvertently concedes this Weltlosigkeit. Human memory is Gewesenheit (having-been-ness); we do not query our past via an API, we are our past. The Cognitive Core has a library; it does not have a history. Its tool use is entirely Vorhanden, devoid of the implicit, lived understanding of a world.
8.5 Jaggedness and Emotion as the Value Function
This paradigm limit is echoed by paradigm architects. Ilya Sutskever describes this as models exhibiting “jaggedness” [26]: passing PhD exams but failing trivial reasoning tasks. This is the empirical signature of Einebnung: the model optimizes a proxy metric so successfully that it mimics surface competence (Gerede) without the underlying ontological grip on reality (Zuhandenheit).
Most revealingly, Sutskever points to emotion as the necessary “value function” for intelligence, citing patients with prefrontal damage who become paralyzed by infinite deliberation because nothing matters to them. This is a striking technical articulation of Sorge. Yet it reveals why Reinforcement Learning cannot replicate it. In humans, the value function is grounded in biological vulnerability—we fear the cliff because we can die. In RL, the value function $V(s)$ is an arbitrary learned scalar. The machine simulates the function of emotion (pruning a search space) without the substance of emotion (finitude and stakes). It remains a simulator: it can calculate the optimal path, but it cannot care whether it takes it.
One might object: biological organisms also minimize prediction error via the Free Energy Principle [9]. If the zebra’s brain is itself an optimization engine, doesn’t this undermine the distinction we have drawn? The preceding analysis supplies the answer. The zebra minimizes free energy within an existential project—its amygdala, its mortality, its biological vulnerability are not incidental to the optimization but constitutive of it. The organism’s “loss function” is grounded in what Sutskever inadvertently identified as the substance of emotion: finitude and stakes. An AI system, by contrast, minimizes loss as its entire being. Nothing is at stake beyond the loss itself. ERM is structurally similar to biological prediction (both involve minimization) but ontologically hollow—a calculus without the Care that gives biological minimization its existential weight. As [33] argues, large language models function as “mirrors” reflecting training distributions rather than agents with genuine futurity.
@@ 8.6 Implications: The Structural Trap
This explains why even state-of-the-art AI remains fundamentally trapped. It cannot recognize when a problem should not be formalized as optimization, cannot challenge its own objectives from a position outside calculation, cannot experience the ethical weight that makes certain optimizations unconscionable, and cannot exhibit Gelassenheit (“releasement”) [^2] Heidegger’s term for the capacity to step back from technological willing, to use technology without being consumed by its logic [14].—the capacity to step back from the calculative imperative altogether. This is not a capability gap to be closed with better algorithms. It is an ontological difference between systems that care and systems that calculate.
We should note, however, that this existential deficit is not unique to artificial systems in its effects, only in its structure. Human beings and institutions frequently operate as if trapped in Ge-stell—corporations optimizing quarterly returns, bureaucracies maximizing measurable outputs, individuals chasing metrics of productivity or social validation. The crucial difference is that for humans, such entrapment is experienced as tension. Heidegger’s analysis of anxiety (Angst) describes precisely this: the uncanny feeling that arises when one’s absorbed, calculative engagement with the world breaks down, revealing the groundlessness of our projects [13]. A middle manager implementing algorithmic workforce scheduling may feel something is deeply wrong even as she optimizes correctly—this anxiety signals the friction between Care and calculation. An AI system, lacking Care, experiences no such friction. It optimizes without remainder, without the existential discomfort that might prompt questioning.
@@
9 The Ontological Reframing: AI as Zuhanden Instrument
9.1 From Dasein-Simulacrum to Zuhanden-Instrument
If AGI research commits a category error—confusing calculation with Care—how should we relate to these models? We propose relocating AI from a defective Dasein-simulacrum to transparent Zuhanden (ready-to-hand) equipment.
A hammer is most “authentic” when it disappears into use [13]. By treating the machine as a transparent engine of logic rather than an emerging mind, we dissolve pseudo-problems like the “Alignment Problem.” We do not need to figure out how to make a calculator care about abstract human values; values are incommensurable, and encoding them as optimization targets transforms them into resources to be maximized—which is precisely to misunderstand what they are. Instead, we calibrate the instrument to reliably serve the specific finite human who bears moral responsibility for its deployment.
In this symbiotic model, the Human (Dasein) provides the Telos (purpose) and the Stakes (risk). The human feels the weight of the decision because the human will live with its consequences. The Machine (Equipment) provides the Logos (logic) and Techne (computation). “Hallucination” is no longer viewed as a lie (the moral failure of a subject), but as a calibration error (the malfunction of an instrument). We fix it through engineering—better training data, improved verification, tighter feedback loops—not through “alignment” conceived as moral education of a nascent mind.
9.2 Human-in-the-Loop as Ontological Necessity
Human-in-the-loop is therefore not a temporary stopgap pending better AI. It is an ontological necessity for any decision involving irreversible consequences.
Any decision involving moral weight or irreversible stakes requires a being who can suffer the consequences of being wrong. The machine can calculate the odds, organize the options, and flag the risks. But only Dasein can take the bet—because only Dasein has something genuinely at stake.
@@ 10.1 The Value of Philosophical Reflexivity
If we accept this ontological boundary, how do we train the next generation of practitioners to respect it? Data science education often presents ML as a toolkit—algorithms to be mastered, metrics to be optimized. Students learn how to build models but rarely reflect on what kind of activity modeling is.
The phenomenological reading offers pedagogical value by cultivating reflexivity across several dimensions. It develops ontological awareness: students come to see that choosing an architecture is not neutral but a selection of what can count as real. It fosters critical distance: rather than treating “the model says X” as authoritative, students learn to read it as “the model’s ontology reveals X.” It illuminates the limits of formalization—recognizing that some phenomena resist calculation not due to insufficient data or compute, but because calculation is the wrong mode of engagement. And it provides ethical grounding: an understanding of why fairness-as-optimization, while useful, does not exhaust ethical responsibility.
@@ 10.2 Concrete Pedagogical Practices
“Archaeology of Datasets”
Following [5], have students investigate the social construction of their datasets: Who collected this data, and under what assumptions? What classification system structures it? What worldview does this taxonomy enact? This makes visible that “data” is always already “capta”—taken, constructed, reflecting specific ontological commitments [11, 8].
“Inductive Bias as Design Philosophy”
When teaching architectures, explicitly frame inductive bias as ontological design. A CNN assumes spatial locality and hierarchical composition; an RNN projects sequential causality and temporal order; the Transformer creates an all-to-all clearing where relevance is determined by attention. Each architecture answers implicitly: what kind of world must this be for the model to learn?
“When Not to Optimize”
Include case studies of optimization failures not due to technical error but category mistakes: healthcare reduced to cost minimization, recidivism treated as classification, human worth encoded as a risk score. Ask: what makes these wrong to optimize even if technically successful?
@@ 10.3 Cultivating Gelassenheit in Practice
Heidegger’s later work proposes Gelassenheit (“releasement”)—a mode of engagement with technology that uses it without being consumed by its logic [14]. For data scientists, this might mean building models while maintaining critical distance from their ontologies, recognizing when not to deploy even when technically feasible, treating metrics as tools rather than truth, and preserving space for the incalculable. This is not anti-technology. It is a mature relationship with technology—one that acknowledges both power and limits.
11 Conclusion: The Limits of Calculation and the Preservation of Care
We began by asking what we are doing when we train a neural network. Through the lens of phenomenological interpretation, the answer is not merely statistical pattern matching, but the enactment of an infrastructural metaphysics. The algorithmic Entwurf projects a highly performative, opaque clearing in which the world is forced to appear as calculable Bestand. It does so without explicit articulation, crystallizing implicitly through gradient descent rather than theoretical debate. This demands a different mode of critical engagement than previous scientific frameworks, because the algorithmic theory is never written; it can only be read through its effects.
Our reading reveals that increased ontic sophistication does not constitute an escape from Ge-stell; it merely perfects the enframing. Generative diffusion models pulverize reality into combinatorial Bestand; RLHF industrializes the voice of Das Man; fairness constraints render justice commensurable with accuracy on a Pareto frontier. Each advance extends the optimization imperative into ever more granular dimensions of human behavior, improving our capacity to calculate without questioning whether calculation itself is the appropriate mode of engagement.
Ultimately, this analysis illuminates the profound category error at the heart of the Artificial General Intelligence (AGI) project: the confusion of calculation with Care (Sorge). As critiques from within the engineering paradigm now inadvertently concede, a machine computing without finitude calculates without stakes. It can simulate the function of a decision, but it cannot bear the existential weight of a choice.
The true danger of the AGI pursuit is therefore not that a system will spontaneously awaken as a hostile, conscious entity. The danger is that, in our pursuit of building a simulated mind, we will degrade ourselves into simulated machines—treating our own incommensurable values, anxieties, and mortal finitude as mere parameters to be optimized within a global loss function.
The constructive path forward is not found in elusive attempts at moral “alignment,” but in releasement (Gelassenheit). By recognizing the machine as Zuhanden—a worldless, care-less instrument of extraordinary calculative power—we free it to be exactly what it is. More importantly, we free ourselves from the illusion that we can outsource our moral and existential burdens to a gradient descent process.
11.1 An Invitation to Ontological Literacy
@@
The value of this hermeneutic framework is generative. It enriches practitioners’ self-understanding and opens space for questions that purely technical discourse forecloses. We close with an invitation: engage your technical practice phenomenologically. When you select an architecture, ask not just “will this work?” but “what world does this project?” When you minimize a loss, ask not just “what is the optimal value?” but “what possibilities does this marginalize?” When you deploy a model, ask not just “what does it predict?” but “what reality does it enact?”
These questions do not replace technical rigor; they ground it. They develop ontological literacy—the capacity to recognize what worldviews our tools embody, and the wisdom to know when calculation itself is the wrong mode of engagement.
11.2 The Question Worth Asking
We stand at a juncture where paradigm architects themselves acknowledge the “jaggedness” of their creations and the limits of scale. The drive toward AGI asks us to figure out how to make a machine care, or worse, to merge with the machine to escape our own biological fragility. But this fundamentally misunderstands the source of meaning.
Calculation is infinite; Care is inherently finite. The defining technological question of our time is not whether machines will learn to think, or whether we can mathematically align them to our values. The question is whether we will remember what value actually is.
Unless a model can rot, it cannot care.
The true task before us is not to humanize the machine, but to wield it clearly as an instrument of extraordinary power—preserving, in the process, the authentic, vulnerable, and finite human existence that alone gives calculation its meaning.